OpenStudy (anonymous):

Similar problem, but still confused... The formula A=P(1+r)^2 gives the amt A in dollars, that P dollars will grow to in 2 years at interest rated r (where r is given as a decimal), using the compound interest? What interest rate will cause 2000 to grow into 2121.80 in two years

OpenStudy (anonymous):

r = sqrt ( A/P) - 1 = sqrt ( 2121.80 / 2000) - 1 = .03 = 3%